### 17 Mar 2026 v1

@@ 2 Algebraic formulation
In this section, we present the mathematical results that provide the foundation for the construction of our algorithms. Our formulation is based on spinorial representations of $\mathrm{Spin}(2n)$. The following lemma shows how to construct orthogonal quantum states in the context of representations of Clifford algebras.
@@
where $\Gamma_{*}$ acts on the $k$-th factor. These operators commute and are self-adjoint. For $\varepsilon_{1},\ldots,\varepsilon_{m}\in\{+,-\}$, define
@@
The following diagrams encode the fact that the chiral decomposition of the tensor product spinor space is compatible with the algebraic action of the Clifford algebra. By means of the canonical projectors onto the positive and negative chiral subspaces, the diagrams illustrate that the action of the total chirality operator intertwines with the projection onto each eigenspace. In particular, it shows that applying the chirality operator before projecting onto a given chiral sector, or first projecting and then applying the restricted action, leads to the same result up to the corresponding eigenvalue.
@@ 3 Spinorial quantum algorithms for classification and search
Proof. Fix an input state $|\psi_{\theta}\rangle\in\mathcal{D}$. A single measurement of the observable $\Gamma_{j}$ produces a random variable
@@
Fix a confidence level $\delta\in(0,1)$. Choosing
@@
In Figure 2, the four histograms obtained for Algorithm 2 illustrate the application of the proposed quantum search protocol to a non-uniform initial distribution encoded in spinorial representations. In all cases, the probability distributions are strongly concentrated around the marked state, indicating that the algorithm successfully amplifies the probability of the target element while suppressing the states that are not solutions.
Figure 2: Probability distributions of Algorithm 2 for two qubits. One observes amplitude amplification in the solution subspace and suppression by destructive interference in the orthogonal complement. Solid bars represent experimental data and hatched bars represent theoretical predictions, showing the high fidelity of Clifford transformations in the segmentation of the state space.
@@ 4 Conclusions and Perspectives
In this work, we developed an algebraic framework based on Clifford algebras and their spinorial representations, which provides a unified language for the construction of quantum algorithms. Within this formalism, we proposed both classification and search algorithms whose functioning is entirely encoded in algebraic operations on spinorial states. One of the main advantages of this approach is that it avoids ad hoc constructions and instead exploits intrinsic algebraic properties, such as orthogonality and anticommutation relations, naturally arising in Clifford algebras.
@@
With respect to the classification algorithm, our proposal allows an arbitrary number of classes. The algorithm is conceptually simple and relies on the construction of mutually orthogonal quantum states, which are naturally obtained from spinorial representations. This orthogonality ensures unambiguous discrimination between classes upon measurement, while the tensor structure guarantees scalability and systematic generalization to higher-dimensional systems.
@@
Concerning quantum search, we analyzed a scenario involving non-uniform initial distribution in which the stored quantum states admit a Clifford encoding, also derived from spinorial representations. In this setting, the oracle implementation becomes particularly simple, as it can be realized using a single generator of the Clifford algebra $Cl(2n)$. This represents a significant simplification when compared to more general oracle constructions, and highlights the computational advantages of encoding information directly in algebraic degrees of freedom.
@@
In a closely related approach, we may also show that a quantum search algorithm based on Hamiltonian simulation can be formulated within the same algebraic framework. By employing spinorial representations and Hamiltonians constructed from generators of the Clifford algebra, the search process can be interpreted as a controlled algebraic rotation in Hilbert space. In this context, the objective is to drive the system from an initial state $|\psi\rangle$ to a target state of the form $\Gamma_{i}|\psi\rangle$. Choosing the Hamiltonian as $H=-i\theta\Gamma_{i}\Gamma_{j}$, the time evolution yields $|\psi^{\prime}\rangle=\exp(-itH)|\psi\rangle=\exp(it\theta\Gamma_{i}\Gamma_{j})\left(\cos(\theta)|\Gamma_{j}\rangle-\sin(\theta)\Gamma_{i}|\Gamma_{j}\rangle\right)=\cos\left[\theta(1+t)\right]|\Gamma_{j}\rangle+\sin\left[\theta(1+t)\right]\Gamma_{i}|\Gamma_{j}\rangle$. An observation at the specific time $t=\left(\pi/2\theta-1\right)$ then results in the state $\Gamma_{i}|\Gamma_{j}\rangle$.
@@
As perspectives, we intend to simulate the proposed classification and search algorithms for systems involving more than two qubits, with the goal of analyzing their scalability, robustness against noise, and potential advantages over standard quantum algorithms. More broadly, these results suggest that Clifford algebras and spinorial methods constitute a promising and mathematically natural framework for the design of quantum algorithms, potentially opening new connections between quantum computation theory, Clifford algebras and spinors.
@@ 5 Appendix
Let $\Gamma_{i}$ and $\Gamma_{j}$ be generators of a Clifford algebra defined in Equation 1. The operator
@@
This provides representation-independent definition of chirality entirely within the framework of Clifford algebras.

### 28 Mar 2026 v2

@@ 2 Algebraic formulation
In this section, we present the mathematical results that provide the basis for the construction of our algorithms. Our formulation is based on spinorial representations of $\mathrm{Spin}(2n)$. The following lemma shows how to construct orthogonal quantum states in the context of representations of Clifford algebras.
@@
where $\Gamma_{*}$ acts on the $k$-th factor. These operators commute and are self-adjoint. For $\varepsilon_{1},\ldots,\varepsilon_{m}\in\{+,-\}$, we have that
@@
The following diagrams illustrate the fact that the chiral decomposition of the tensor product spinor space is compatible with the algebraic action of the Clifford algebra. By means of the canonical projectors onto the positive and negative chiral subspaces, the diagrams illustrate that the action of the total chirality operator intertwines with the projection onto each eigenspace. In particular, it shows that applying the chirality operator before projecting onto a given chiral sector, or first projecting and then applying the restricted action, leads to the same result up to the corresponding eigenvalue.
@@ 3 Spinorial quantum algorithms for classification and search
Proof. Let $|\psi_{\theta}\rangle\in\mathcal{D}$ be an inpute state. A single measurement of the observable $\Gamma_{j}$ produces a random variable
@@
Let $\delta\in(0,1)$ be a confidence level. Choosing
@@
In Figure 2, the four histograms obtained for Algorithm 2 illustrate the application of the proposed quantum search protocol to a non-uniform initial distribution based on spinorial representations. In all cases, the probability distributions are strongly concentrated around the marked state, indicating that the algorithm successfully amplifies the probability of the target element while suppressing the states that are not solutions.
Figure 2: Probability distributions of Algorithm 2 for two 2,3,4 and 5 qubits. One observes amplitude amplification in the solution subspace and suppression by destructive interference in the orthogonal complement. Solid bars represent experimental data and hatched bars represent theoretical predictions, showing the high fidelity of Clifford transformations in the decomposition of the state space.
@@ 4 Conclusions and Perspectives
In this work, we developed an algebraic formulation based on Clifford algebras and their spinorial representations, which provides a unified language for the construction of quantum algorithms. Within this formalism, we proposed both classification and search algorithms whose functioning is entirely based on algebraic operations on spinorial states. One of the main advantages of this approach is that it avoids ad hoc constructions and instead exploits intrinsic algebraic properties, such as orthogonality and anticommutation relations, arising in Clifford algebras.
@@
With respect to the classification algorithm, our proposal allows an arbitrary number of classes. The algorithm is conceptually simple and relies on the construction of mutually orthogonal quantum states, which are obtained from spinorial representations. This orthogonality ensures unambiguous discrimination between classes upon measurement, while the tensor structure guarantees systematic generalization to higher-dimensional systems.
@@
Concerning quantum search, we analyzed a scenario involving non-uniform initial distribution in which the stored quantum states admit a Clifford encoding, also derived from spinorial representations. In this scenario, the oracle implementation becomes particularly simple, as it can be realized using a single generator of the Clifford algebra $Cl(2n)$. This represents a significant simplification when compared to more general oracle constructions, and highlights the computational advantages of encoding information directly in algebraic structures.
@@
In a closely related approach, we may also show that a quantum search algorithm based on Hamiltonian simulation can be formulated within the same algebraic structure. By employing spinorial representations and Hamiltonians constructed from generators of the Clifford algebra, the search process can be interpreted as a controlled algebraic rotation in Hilbert space. In this context, the objective is to drive the system from an initial state $|\psi\rangle$ to a target state of the form $\Gamma_{i}|\psi\rangle$. Choosing the Hamiltonian as $H=-i\theta\Gamma_{i}\Gamma_{j}$, the time evolution yields $|\psi^{\prime}\rangle=\exp(-itH)|\psi\rangle=\exp(it\theta\Gamma_{i}\Gamma_{j})\left(\cos(\theta)|\Gamma_{j}\rangle-\sin(\theta)\Gamma_{i}|\Gamma_{j}\rangle\right)=\cos\left[\theta(1+t)\right]|\Gamma_{j}\rangle+\sin\left[\theta(1+t)\right]\Gamma_{i}|\Gamma_{j}\rangle$. An observation at the specific time $t=\left(\pi/2\theta-1\right)$ then results in the state $\Gamma_{i}|\Gamma_{j}\rangle$.
@@
More broadly, these results suggest that Clifford algebras and spinorial methods constitute a promising formulation for the design of quantum algorithms, potentially opening new connections between quantum computation theory, Clifford algebras and spinors.
As perspectives, we intend to simulate the proposed classification and search algorithms for systems involving more than five qubits on better-calibrated quantum devices, with the goal of analyzing their scalability, robustness against noise, and potential advantages over standard quantum algorithms.
@@ 5 Appendix
Let $\Gamma_{i}$ and $\Gamma_{j}$ be generators of a Clifford algebra defined in the eq.(1). The operator
@@
This provides representation-independent definition of chirality entirely within the structure of Clifford algebras.